# Horizon Activation Mapping for Neural Networks in Time Series Forecasting

### v2

@@ 1 Introduction
Time series forecasting has seen performance improvements because of neural network models similar to those in other modalities in the recent years and these models have been able to forecast over longer horizons. The existence of many families of neural network models that perform well in the forecasting task has motivated arguments for and against models of different families for the task. The arguments address the speed and memory aspects of the forecasting, with performances often remaining similar [11, 13]. Foundation models have also been introduced towards the task that are capable of zero-shot generalizability subject to some limitations [3, 5]. The model families have yet to find large enough supervised learning datasets towards comparisons across them and this motivates a framework that can analyze gradient activities in these networks agnostic to the types of layers in them.
@@ 3.2 Batch Size Differences
Even though there is variability in magnitudes across batch sizes in Figures 2(c) and 4, Figure 3 shows a surprising polynomial curve across batch sizes in converged models over a single epoch, motivating future work.
@@ 3.3 Early Stopping
Figure 5(d) shows the gradient norm averages reducing in magnitude as the number of epochs after early stopping increases, indicating that increase in errors correspond to magnitude reductions in early stopping. From the HAM and area curves, the causal mode curves reduce in magnitude after early stopping, first steeply and then gradually, with the magnitude reductions corresponding to being in a local minima. Similarly, all the curves in the difference plot are below the uniform line and that of the early stopped model is closest to it throughout the horizon. The gradient equivariant point of the early stopped model is closer to $\frac{H}{2}$ than the other models’ suggesting that activities in longer regions of the horizon don’t improve the performance.
Epochs MSE MAE
20 0.224 0.341
30 0.227 0.345
40 0.229 0.347
50 0.233 0.350
@@ 4.2 Difference Plots across Model Families
Figure 9: Difference plot between gradient norm averages of the first $h$ and last HhH-h timesteps in a horizon of size $H$. The points below the curves are gradient equivariant points of the models.
As a general rule, curves greater than uniform differences focus on shorter horizons more and vice versa, and over corresponding subseries lengths with respect to gradient equivariant points in Figure 9. Pyraformer’s difference curves address long horizon forecasting as an activity magnitude-based emphasis of shorter horizons and correspond to its largest errors across all models. Multi-Resolution DDPM’s curves decrease more after its gradient equivariant points in an increasing order of horizon size with curves before them remaining more than proportionate. Its equivariant points are all closest to the $\frac{H}{2}$ timesteps highlighting that they are able to use noise distributions to address long horizon forecasting in terms of the norm average magnitudes. FEDformer’s difference plots decrease in magnitudes after equivariant points between H=96,192H={96,192} and H=336,720H={336,720} models indicating the robustness of its self-attention in adapting to data distributions and decreases corresponding to longer regions in the horizon as necessary. SpaceTime’s difference curves emphasize increases in shorter regions of the horizon decreasing in $H=\{96,192\}$ models and along with its gradient equivariant points being farthest away in the horizons, use its weights’ power distributions to exponentially highlight longer regions of the horizon using relative activity magnitudes. NHITS’ difference plots resemble SpaceTime’s in its H=96,192H={96,192} models and show reductions in its anti-causal mode’s longer subseries more than its causal mode’s when horizon size increases from 336 to 720. The difference plot decreases in magnitude as the causal mode includes the last 18% and 14% of the timesteps in $H=336$ and $H=720$ models respectively, favouring longer horizons lesser as horizon sizes increase.

### v3

@@ 1 Introduction
Time series forecasting has seen performance improvements because of neural network models similar to those in other modalities in the recent years and these models have been able to forecast over longer horizons. The existence of many families of neural network models that perform well in the forecasting task has motivated arguments for and against models of different families for the task. The arguments address the speed and memory aspects of the forecasting, with performances often remaining similar [12, 14]. Foundation models have also been introduced towards the task that are capable of zero-shot generalizability subject to some limitations [3, 5]. The model families have yet to find a large enough supervised learning dataset towards comparisons across them and this motivates a framework that can analyze gradient activities in these networks agnostic to the types of layers in them.
@@ 3.2 Batch Size Differences
Even though there is variability in magnitudes across batch sizes in Figures 2(c) and 3, Figure 4 shows a surprising polynomial curve across batch sizes in converged models over a single epoch, motivating future work.
@@ 3.3 Early Stopping
Figure 5(c) shows the gradient norm averages reducing in magnitude as the number of epochs after early stopping increases, indicating that increase in errors correspond to magnitude reductions in early stopping. From the HAM and area curves, the causal mode curves reduce in magnitude after early stopping, first steeply and then gradually, with the magnitude reductions corresponding to being in a local minima. Similarly, all the curves in the difference plot are below the uniform line and that of the early stopped model is closest to it throughout the horizon. The gradient equivariant point of the early stopped model is closer to $\frac{H}{2}$ than the other models’ suggesting that activities in longer regions of the horizon don’t improve the performance.
@@ 4.2 Difference Plots across Model Families
Figure 9: Difference plots between gradient norm averages of first $h$ and last HhH-h timesteps in a horizon of size $H$ with gradient equivariant points below them.
Pyraformer’s difference curves address long horizon forecasting as an activity magnitude-based emphasis of shorter horizons and correspond to its largest errors across all models. Multi-Resolution DDPM’s curves decrease more after its gradient equivariant points in an increasing order of horizon size with curves before them remaining more than proportionate. Its equivariant points are all closest to the $\frac{H}{2}$ timesteps highlighting that they are able to use noise distributions to address long horizon forecasting in terms of the norm average magnitudes. FEDformer’s difference plots decrease in magnitudes after equivariant points between H=96,192H={96,192} and H=336,720H={336,720} models indicating the robustness of its self-attention in adapting to data distributions and decreases corresponding to longer regions in the horizon as necessary. SpaceTime’s difference curves emphasize increases in shorter regions of the horizon decreasing in $H=\{96,192\}$ models and along with its gradient equivariant points being farthest away in the horizons, use its weights’ power distributions to exponentially highlight longer regions of the horizon using relative activity magnitudes. NHITS’ difference plots resemble SpaceTime’s in its H=96,192H={96,192} models and show reductions in its anti-causal mode’s longer subseries more than its causal mode’s when horizon size increases from 336 to 720. The difference plot decreases in magnitude as the causal mode includes the last 18% and 14% of the timesteps in $H=336$ and $H=720$ models respectively, favouring longer horizons lesser as horizon sizes increase.